# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@
Abstract.
Understanding how capabilities evolve into core capabilities—and how core capabilities may ossify into rigidities—is central to innovation strategy (Leonard-Barton 1992, Teece 2009). A major challenge in formalizing this process lies in the interactive nature of innovation: successes in one domain often reshape others, endogenizing specialization and complicating isolated modeling. This is especially true in ecosystems where firm capabilities and innovation outcomes hinge on managing interdependencies and complementarities (Jacobides et al. 2018, 2024).
@@ Abstract.
Key-words: Innovation Ecosystems, Dynamic Capabilities, Interacting Bernoulli Processes, Patent Interactions, Technological Specialization
1. Introduction
Understanding the extent to which capabilities evolve into core capabilities—and how core capabilities may turn into core rigidities (Leonard-Barton 1992)—is a crucial question for firms, particularly those that are innovation-driven (Teece 2009, Eisenhardt and Martin 2000, Helfat and Peteraf 2003). Salter and Alexy (2014) support this view by emphasizing that a well-balanced mix of skills within teams is a fundamental driver of innovation—particularly when these skills are leveraged within a proactive ecosystem capable of transforming them into core capabilities (Jacobides et al. 2018). Finding such a mix of skills within an ecosystem is, however not an easy task and further exploration is needed. An interesting hint in this direction is provided in Kim and Magee (2017) and in Pichler et al. (2020). They demonstrate that knowledge flows predominantly occur within previously linked technological domains, which implies that domain-specific learning can reinforce the development of specialized skill sets (Cohen et al. 1990, Malerba and Orsenigo 1997). In their study, as well as in several others (Teece et al. 1997, Adner and Kapoor 2010, Youn et al. 2015, Pugliese et al. 2019) it clearly emerges how the concept of skills and capabilities is inherently linked with the concept of technological fields. Studying how technologies, inherently linked to specific skill requirements, interact within an ecosystem is, therefore, crucial for formalizing the transformation of these skills into core capabilities (Granstrand 1998, Kim et al. 2022). One of the main challenges posed by the interaction among innovations is that they endogenize the processes of specialization and the formation of core capabilities. In other words, the dynamic and path-dependent nature of capability evolution (Teece 2009, Eisenhardt and Martin 2000) is shaped by the co-evolution of technologies and markets. As a result, the interplay among innovations does more than merely shifting the external environment—it recursively shapes the internal structure and evolutionary trajectory of firm capabilities, making it difficult to model them in isolation. This challenge is well recognized in the literature, particularly given that firm capabilities operate within ecosystems whose value creation depends on how multilateral interdependencies–based on various forms of complementarities–are managed (Jacobides et al. 2018). Both generic and idiosyncratic complementarities across capabilities– and thus across technological domains– are crucial determinants of a firm’s success in innovation (Jacobides et al. 2024). An example of this is given by innovation-based platforms, where developers must be incentivized to generate complementary innovations (Jacobides et al. 2024).
Given the above premises, formalizing the processes that govern interactions among innovations across different technological domains becomes essential for determining the extent to which ecosystems can effectively exploit complementarities—thereby fostering the development of core capabilities while avoiding their degeneration into core rigidities. The aim of the present work is to develop a novel mathematical framework to rigorously formalize such processes of interactive innovations.
Technological innovation is not merely a process of isolated breakthroughs but is strongly influenced by self-reinforcing structures, where mutually connected technological fields benefit from each other, leading to cumulative innovation dynamics (Salter and Alexy 2014, Youn et al. 2015, Napolitano et al. 2018, Castaldi et al. 2015). The relationship between cross-category interactions and innovation success is, however, not necessarily straightforward and necessitates a dedicated formal treatment. While interactions within closely related Cooperative Patent Classification (CPC) categories may increase the overall likelihood of incremental innovation, interactions across technological domains may be a key driver of breakthrough innovations—though such events remain rare (Castaldi et al. 2015). Kovács et al. (2021), for instance, show that patents classified in high-contrast categories—those with clear boundaries from neighboring technologies—receive more attention and citations than those in low-contrast categories. This suggests that classification is not merely an administrative process but actively shapes innovation diffusion by influencing how patents are discovered, cited, and integrated into future technological developments. According to Kovács et al. (2021), one key mechanism through which this occurs is search dynamics—the way patent users identify and engage with prior knowledge. Their findings reinforce the idea that the way technologies are classified and interconnected affects not only their immediate visibility but also their long-term influence on the technological evolution of firms, seen as innovation ecosystems (Coccia 2017, Colladon et al. 2025). Specifically, understanding these systemic interdependencies in technologies is central to studying the logic of search dynamics and, consequently, to modeling how capabilities evolve within firms through a clearer understanding of their technological evolution (Pichler et al. 2020).
The literature on this topic has primarily documented the clustering of technological domains and their reinforcing interactions (Youn et al. 2015), yet it often falls short in explicating the underlying mechanisms—highlighting the need for a formal model. Indeed, while Leonard-Barton (1992) emphasized that core capabilities may evolve into core rigidities if not continuously adapted, numerous scholars over the years have argued that a formal framework is indispensable for meaningfully quantifying and understanding this transformation (Teece 2009, Zollo and Winter 2002, Helfat and Peteraf 2003, Youn et al. 2015). The latter should model the recursive, path-dependent interactions among innovations and pave the way for future research to assess how these interactions influence the development, entrenchment, or renewal of capabilities. Our study bridges this gap by introducing a mathematical framework to assess how innovation success in one CPC category reinforces success, not only in the same category, but also in the other categories, thus offering a structured explanation for how technological trajectories emerge and evolve. A major difficulty in building such a formal framework lies in the fact that, despite the growing recognition of cross-category knowledge flows (Blazsek and Escribano 2010, Nemet 2012, Pichler et al. 2020, Singh et al. 2021), the role of cross-category interactions in innovation remains largely unexplored in formal models.
The relevance of our model and its ability to explain real-world innovation dynamics is assessed by identifying a set of stylized facts some of which have been consistently documented in the empirical literature. While existing models typically focus on a subset of these phenomena, our framework provides a unified theoretical explanation that links them through the structure of interactions among technological domains as required by the literature (Pichler et al. 2020). In few words, our model consists of a finite set of interacting Bernoulli processes, each for each technological category. The time-steps are marked by the succession of patents, according to their full date of publications. The status $1$ of the process related to category $h$ identifies a patent which is a “success”, with respect to a novelty measure, in the technological field $h$. The dynamics of the Bernoulli processes are governed by a self- and cross-category reinforcement rule: the probability that patent $n+1$ will be a success for category $h$ has an increasing dependence on the number of past successes, not only in category $h$, but also in all the other categories. The interaction matrix $\Gamma=(\gamma_{j,h})_{j,h}$ collects the weight of this dependence: $\gamma_{j,h}\geq 0$ quantifies how much past successes in a given category $j$ induces future successes in category $h$. As a novelty measure for our model, we use a variant of the forward citation index in Squicciarini et al. (2013): essentially, in order to measure the impact of a given patent in a certain CPC-1 category, we measure the number of citations to the considered patent by subsequent patents belonging to the considered category, normalized with respect to the patent cohort. However, it is very important to note that our methodology is presented in general so that it can be applied with different classifications in categories and/or different novelty measures, even in other contexts. [^1] The interested reader is referred to Section A in appendix for a discussion on novelty measures. There we also motivate our preference for the index of Squicciarini et al. (2013) over other measures of patent novelty.
All the provided theoretical results are proven rigorously and are supported by the analysis of a real-world data set from the GLOBAL PATSTAT database. More specifically, we analyze all the patents published in the period $[1980-2018]$ showing that the observed behaviors match with the theoretical ones. We also perform a statistical study on the interaction intensity among the categories.
Below, we present a list of core empirical regularities that our model is able to jointly explain, offering a more comprehensive perspective on the nature of technological path dependence. We divided the mentioned regularities into those related to the formation of core capabilities—denoted as $\mathcal{C}$.i—and those related to the ossification of core capabilities into core rigidities—denoted as $\mathcal{R}$.i.
  • $\mathcal{C}$1
    Early-stage interventions produce stronger and more persistent effects, especially before the innovation landscape fragments into relatively autonomous and specialized domains. [^2] Similar calls for timely and anticipatory policy actions have been raised in studies emphasizing the cumulative and path-dependent nature of technological change (David 1985, Mazzucato 2018). Our model is ruled by a self- and cross- reinforcement mechanism so that the probability of a future success in a certain category depends on the number of past successes in that category and also in all the other categories. This study therefore highlights the importance of early interventions as having the strongest systemic impact, thereby reinforcing Leonard-Barton (1992)’s view that the evolution of core capabilities is path-dependent and resistant to late-stage adjustments—emphasizing the strategic imperative for timely, anticipatory actions in innovation policy and capability building.
  • $\mathcal{C}$2
    The probability of observing a future success decreases along time. This fact is reflected in a sub-linear growth of the number of observed successes: indeed, for each category $h$, the number $S_{t,h}$ of successes in category $h$ exhibits a power-law growth with an Heaps’ exponent lower than one. In other words, technological categories become ”crowded” over time, with increasing saturation and depletion of easy innovation opportunities, making new breakthroughs less frequent (Jones 2009, Bloom et al. 2020, Clancy 2023). In a broader innovation ecosystem, this supports the need for cross-category recombination: when a single field is exhausted, future breakthroughs often require combining insights from other fields (Fleming 2001), thereby creating new opportunities of developing core capabilities.
  • $\mathcal{C}$3
    Irreducibility of the interaction matrix: all the categories present the same Heaps exponent and this means that all the categories interact each other or directly (when $\gamma_{j,h}>0$) or undirectly by means of a path in the graph associated to the interaction matrix $\Gamma$. This evidence formalizes the arguments in Weitzman (1998), Fleming (2001), Alves et al. (2007), Godin (2017) among others.
  • $\mathcal{R}$1
    Relative number of successes across categories stabilizes along time: for two different categories $h$ and $j$, the ratio $S_{t,h}/S_{t,j}$ stabilizes over time toward the ratio of their respective components of the eigenvector-centrality score vector ${\mathbf{u}}$, that is $S_{t,h}/S_{t,j}\to u_{h}/u_{j}$. This means that the relative proportions of successes across categories remain stable in the long run. Such stability reflects the consolidation of core capabilities within firms constituting the innovation ecosystem, on the one hand paving the way for the formation of new core capabilities and on the other starting the process of ossification of traditional ones into core rigidities (Leonard-Barton 1992). While related ideas regarding the influence of central nodes over time in economic production networks have been theorized (Acemoglu et al. 2016), to the best of our knowledge, this specific idea has never been formally stated or formalized within the innovation literature. [^3] Although empirical evidence point in this direction (Breschi et al. 2000, Malerba 2002).
  • $\mathcal{R}$2
    The long-run decline of the cross-category correlations: if we denote by $X_{n,h}$ the random variable that takes the value $1$ if patent $n$ is a success in category $h$ and 00 otherwise, then the correlation coefficients between $X_{n,h}$ and $X_{n,j}$ decreases to zero along time. Indeed, although the interaction matrix $\Gamma$ is constant along time, its impact on the probabilities of success decreases linearly along time, generating vanishing correlations among the categories. This last piece of evidence is reminiscent of the knowledge spillovers literature (Jaffe and Trajtenberg 2002). In their book (and subsequent works), Jaffe and Trajtenberg (2002) observe increasing specialization within technological fields, implying that cross-field citations — and, by extension, indirect correlations — tend to diminish over time relative to within-field citations. However, the treatment in Jaffe and Trajtenberg (2002) remains primarily empirical and descriptive. This process of specialization is at the basis of the ossification process into core rigidities described in Leonard-Barton (1992).
A more thorough positioning of our contribution within the existing literature, considering the empirical regularities discussed above, is presented in Section 1.1.
Our novel mathematical model offers a strategic tool for policymakers and practitioners. Indeed, as aforementioned, it helps managers understanding how and when innovation processes interact, when a technological field becomes saturated, and how it relates to other fields over the long term. This, in turn, provides valuable insights to identify how and when the capabilities—required for developing innovations within a given technological field—may evolve into core capabilities or degenerate into core rigidities (Leonard-Barton 1992).
Furthermore, our formal model can guide managers in innovation-related decision-making, which has become more challenging due to the growing complexity and variety of innovation ecosystems (Dias Sant´ Ana et al. 2020), as well as the increasing interdependencies among technologies (Adner and Kapoor 2010, Russell and Smorodinskaya 2018, Granstrand and Holgersson 2020) –especially in an environment where innovation ecosystems coevolve (Thomas and Autio 2019, Paasi et al. 2023). In particular, we illustrate how a localized innovation shock in one technological domain can propagate across the broader ecosystem, influencing success rates in other domains depending on the structure of interdependencies. Importantly, we show that even targeted interventions can have far-reaching effects—either amplifying or constraining innovation elsewhere—thereby offering managers additional tools to assess competitive dynamics and make strategic decisions within innovation ecosystems.
In the next subsection we position our work in relation to the existing literature.
1.1. Positioning within the literature
Prior studies in the management literature have generally addressed aspects of patent dynamics—such as citation behavior (Lee et al. 2012), innovation competition (Keller 2007), and R&D investment (Akcigit and Kerr 2018, Aghion and Howitt 2007)— lacking, however, a formalization of the process governing cross-category technological interactions. In the course of the years scholars have deepen the latter aspect by investigating knowledge spillovers across technologies (Kim and Magee 2017) and forming patents’ networks (Higham et al. 2022). Emerging methods, including machine learning models for predicting patent success (Yang 2023), and descriptive studies on radical inventions’ cross-domain influence (Dahlin and Behrens 2005), highlight this gap. These works, however, remain empirical in nature and do not provide a formal understanding of the processes that drive their results. The present paper addresses this by explicitly capturing – via interactive Bernoulli processes– how innovation in one CPC category affects outcomes in others, providing a structural model of technological co-evolution.
Parallel developments in management research, often leveraging NLP and LLMs, have focused on patent similarity across categories (Younge and Kuhn 2016, Arts et al. 2018, Jurek 2024, Bekamiri et al. 2024). While useful for mapping static proximity, these approaches remain correlational and fall short in explaining dynamic mechanisms such as technological convergence, divergence, and the long-run distribution of innovation outcomes (Ganguli et al. 2024). Even studies incorporating dynamics (Kim and Magee 2017, Park and Magee 2017, Taalbi 2023, Maragakis et al. 2023) tend to remain empirical, lacking formal structure.
As innovation increasingly unfolds through technological interdependencies over time – generative AI constitutes an example of this phenomenon (Chen et al. 2023, Du et al. 2024, Mohammadabadi 2025)– recent works emphasize the need for formal models that endogenize these relationships (Pichler et al. 2020, Taalbi 2023). Existing models, in particular, are often capable of showing isolated phenomena but do not jointly address long-run specialization, capability evolution, or structural path dependence. Our contribution lies in bridging this gap by introducing a unified framework that models interacting technological domains, capturing features such as long-run dominance, category cycles, and the evolution from interdisciplinarity to specialization (Wang and Schneider 2020).
We now consider the main features of our model with respect to the literature:
  • Early-stage interventions produce stronger and more persistent effects. Back at the time David (1985) introduced the concept of path dependence in technological change, illustrating how early choices can lock-in trajectories that resist later correction. This inertia can either evolve into rigidities if not shaped proactively (Leonard-Barton 1992), or be catalyzed into core capabilities through well-designed, mission-oriented innovation policies implemented by institutions (Mazzucato 2018, Kattel and Mazzucato 2018). Our paper elucidates these dynamics by examining how innovation processes interact with one another, thereby potentially influencing the early-stage outcomes of adjacent processes.
  • The probability of observing a future success decreases along time. This aspect is reflected in a power-law growth of the number of observed successes with an exponent smaller than one. This evidence was noted in Heaps (1978) (indeed, the power-law exponent is often called the Heaps’ exponent) and observed in many fields, such as economics, natural sciences, linguistics and so on. Since Heaps’ work, it has become natural to require that applicative models satisfy this power-law property (Clauset et al. 2009, Mitzenmacher 2004, and 2005).
  • All categories interact with one another, either directly or indirectly. Several studies in the literature have hypothesized or empirically observed this pattern (Verspagen 2007, Rigby 2015, Higham et al. 2022). Few, however have provided a formal modelization of this phenomenon. We did so by introducing an interaction matrix governing the probability of future success in a certain category.
  • Relative number of successes across categories stabilizes along time. Despite continuous innovation, patent categories exhibit long-term stability in their relative success rates. We demonstrate that the ratio of cumulative successes across categories converges to the ratio of their centrality scores, indicating a predictable long-run distribution. This fact aligns with the empirical observation that technological domains rarely disappear entirely (Park and Magee 2017, Kim and Magee 2017) and an interpretation can be given in light of the work of Maslach (2016): firms persist after incremental failures but pivot after radical ones. As most innovation is incremental in nature, firms tend to remain anchored in familiar technological domains, reinforcing their sustained importance despite the emergence of new fields. [^4] This also aligns with empirical observations that technological domains, such as pharmaceuticals and automotive, maintain consistent importance over time (Maslach 2016, Perri et al. 2020, Lafond and Kim 2019)
  • The long-run decline of the cross-category correlations. Our model captures the declining correlation between patent categories over time, reflecting a trend toward increased specialization –a well-documented phenomenon in the literature (Ganguli et al. 2024). Namely, technologies that were initially interdisciplinary tend to become more specialized as they mature. A well-documented historical example of this phenomenon is the divergence between telecommunications and information technology (IT) (Messerschmitt 1996, Shin 2010). [^5] Initially, these domains were tightly interconnected, with innovations in data transmission and computing co-evolving (Jones and McGuirt 1991, Fransman 2001, Leiner et al. 2009). Over time, however, they have become increasingly specialized: telecommunications has focused on network infrastructure and communication protocols, while IT has expanded into computing systems, software, and data management (Messerschmitt 1996, Garrone et al. 2002).

### v2

@@
Abstract
Understanding how capabilities evolve into core capabilities—and how core capabilities may ossify into rigidities—is central to innovation strategy (Leonard-Barton 1992, Teece 2009). A major challenge in formalizing this process lies in the interactive nature of innovation: successes in one domain often reshape others, endogenizing specialization and complicating isolated modeling. This is especially true in ecosystems where firm capabilities and innovation outcomes hinge on managing interdependencies and complementarities (Jacobides, Cennamo and Gawer 2018, 2024).
@@ Abstract
Technological Specialization,
keywords:
1 Introduction
Understanding how capabilities evolve into core capabilities—and how these may ossify into core rigidities (Leonard-Barton 1992)—is central for innovation-driven firms (Teece 2009, Eisenhardt and Martin 2000, Helfat and Peteraf 2003). A balanced portfolio of skills within teams, embedded in an ecosystem that can mobilize complementarities, is a key driver of innovation (Salter and Alexy 2014, Jacobides, Cennamo and Gawer 2018). Yet, assembling and orchestrating such skills is challenging. Evidence shows that knowledge flows occur predominantly within previously linked technological domains (Kim and Magee 2017, Pichler, Lafond and Farmer 2020), reinforcing domain-specific learning (Cohen et al. 1990, Malerba and Orsenigo 1997). Hence, skills and capabilities are inherently tied to whether or not the technological fields of teams’ members interact and in which way they do so (Teece, Pisano and Shuen 1997, Adner and Kapoor 2010, Youn et al. 2015, Pugliese et al. 2019); formalizing how technologies interact within ecosystems is, therefore, essential to understand when capabilities become core—and when they risk becoming rigidities (Granstrand 1998, Kim et al. 2022).
A primary challenge is that interactions among innovations endogenize specialization and capability formation. Capability evolution is path-dependent (Teece 2009, Eisenhardt and Martin 2000) and co-evolves with technologies and markets. Ecosystem value creation depends on managing multilateral interdependencies rooted in complementarities (Jacobides, Cennamo and Gawer 2018), both generic and idiosyncratic, which materially shape innovation outcomes (Jacobides, Cennamo and Gawer 2024)—as in platform settings where complementary innovation must be incentivized.
Aim and approach
Firms and economies do not innovate in isolation: successes in one domain often reshape others, altering trajectories of specialization and creating path-dependent dynamics that resist later correction (Leonard-Barton 1992, Teece 2009). We develop a mathematical framework that formalizes interactive innovation across technological domains. Our model builds on a family of interacting reinforced Bernoulli processes, indexed by CPC categories, in which the probability of a future success in one category depends not only on its own past successes but also on the cumulative successes in other, related categories: at each patent time-step, the process in category $h$ takes value $1$ if the patent is a “success” in $h$ (according to a novelty/impact threshold), and 0 otherwise. Success probabilities evolve via self- and cross-category reinforcement: the probability that patent n+1n{+}1 is a success in $h$ increases with past successes in $h$ and in all other categories, with effects weighted by a nonnegative interaction matrix $\Gamma=(\gamma_{j,h})$. This setup captures cumulative, path-dependent innovation and cross-domain spillovers documented in the literature (Salter and Alexy 2014, Youn et al. 2015, Napolitano et al. 2018, Castaldi, Frenken and Los 2015) and how shocks in one domain propagate across the ecosystem.
As our baseline novelty/impact measure, we adopt a category-level variant of the forward citation index in Squicciarini, Dernis and Criscuolo (2013): for a given patent and CPC-1 category, we count subsequent citations from the same category and normalize by cohort. The modeling framework is general and can accommodate alternative taxonomies and novelty definitions. [^1] See Appendix A for a discussion of novelty measures and the rationale for our baseline choice.
Why a formal model is needed
While relatedness and clustering of technologies are well documented (Youn et al. 2015), the mechanisms through which cross-category interactions shape innovation trajectories remain under-theorized. Classification affects diffusion and attention—patents in high-contrast categories draw more citations (Kovács, Carnabuci and Wezel 2021)—implying that the structure of technological interdependencies governs search and long-run influence (Coccia 2017, Colladon, Guardabascio and Venturini 2025, Pichler, Lafond and Farmer 2020). A formal model that endogenizes these interactions is thus necessary to quantify how capabilities are formed, entrenched, or renewed.
Empirical scope and validation
We analyze GLOBAL PATSTAT patents (1980–2018), show that observed behaviors align with the model’s predictions, and study the intensity of cross-category interactions. We further examine how localized innovation shocks propagate across the CPC network.
Stylized facts our model jointly explains
We organize the empirical regularities into those linked to the emergence of core capabilities ($\mathcal{C}$.i) and to their ossification into core rigidities ($\mathcal{R}$.i):
  • $\mathcal{C}$.1
    Early interventions are disproportionately potent. Because success probabilities are reinforced by accumulated successes across categories, early shocks/policies have stronger and more persistent ecosystem-wide effects (cf. David 1985, Mazzucato 2018). This formalizes the path dependence emphasized by Leonard-Barton (1992).
  • $\mathcal{C}$.2
    Declining success probabilities over time. Each category’s cumulative successes $S_{t,h}$ grow sublinearly with a Heaps-type power law (exponent <1<1), consistent with crowding/saturation of opportunities (Jones 2009, Bloom et al. 2020, Clancy 2023). This underscores the value of cross-category recombination when single fields mature (Fleming 2001).
  • $\mathcal{C}$.3
    Irreducible interdependence. A common Heaps exponent across categories indicates a connected interaction structure (directly or via paths in the graph of $\Gamma$), aligning with recombinant-search arguments (Weitzman 1998, Fleming 2001, Alves et al. 2007, Godin 2017).
  • $\mathcal{R}$.1
    Stabilization of relative success shares. For categories $h$ and $j$, $S_{t,h}/S_{t,j}\to u_{h}/u_{j}$, where $\mathbf{u}$} is the eigenvector-centrality of the interaction network. Long-run stability reflects consolidation of core capabilities and nascent rigidity (Leonard-Barton 1992); related ideas appear in production networks (Acemoglu, Akcigit and Kerr 2016), with suggestive evidence in innovation (Breschi, Malerba and Orsenigo 2000, Malerba 2002).
  • $\mathcal{R}$.2
    Vanishing cross-category correlations. If $X_{n,h}$ indicates a success in $h$ at patent $n$, $\mathrm{corr}(X_{n,h},X_{n,j})\to 0$ over time as reinforcement effects dilute, echoing specialization dynamics in knowledge spillovers (Jaffe and Trajtenberg 2002).
Relation to the literature
Our work contributes to several strands of the economics literature. First, it formalizes the dynamics of R&D spillovers and network complementarities (Jaffe and Trajtenberg 2002, Kim and Magee 2017, Higham, Contisciani and De Bacco 2022), providing a structural model of how innovation in one Cooperative Patent Classification (CPC) category affects others. Second, it advances the study of path dependence in innovation (David 1985, Mazzucato 2018, Pichler, Lafond and Farmer 2020), showing how early shocks and interventions lock in trajectories of specialization. Third, it builds on the literature on firm capabilities and dynamic evolution (Eisenhardt and Martin 2000, Helfat and Peteraf 2003), highlighting how complementarities across technologies mediate the transformation of capabilities into assets or rigidities. Fourth, it complements recent empirical work using machine learning, NLP, and network methods to map technological proximity (Younge and Kuhn 2016, Arts, Cassiman and Gomez 2018, Jurek 2024, Bekamiri, Hain and Jurowetzki 2024), by providing a formal dynamic mechanism that explains observed patterns of convergence, divergence, and long-run specialization. Finally, it connects to the historical literature on lock-in and recombination (Weitzman 1998, Fleming 2001, Verspagen 2007, Rigby 2015), offering a unified framework that ties together sublinear growth, relative stability, and diminishing cross-field correlations.
Clarification on Novelty, Innovation, and Invention